The electron beam in a colour TV is accelerated through 32 kV and then strikes the screen. What is the wavelength of the most energetic X-ray photon?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
V = 32 KV = 32 * 10^3 V When accelerated through 32 KV λ = hc/E = 1242/32* 10^3 = 38.8 * 10^-3 nm = 38.8 pm.